# The Last Iterate Advantage: Empirical Auditing and Principled Heuristic Analysis of Differentially Private SGD

### v2

@@ A.1 Computing $\delta$ from $\varepsilon$
These expressions can be evaluated efficiently. Since $f$ is monotone, it can be inverted via binary search. We can also use binary search to evaluate $\varepsilon$ as a function of $\delta$.
@@ Appendix B Linear Worst Case for Full Batch Setting
We can formalize this using the language of Gaussian DP [DRS19]: DP-GD (Algorithm 1 with $q=1$) satisfies $\sqrt{T}/\sigma$-GDP. (Each iteration satisfies $1/\sigma$-GDP and adaptive composition implies the overall guarantee.) This means that the privacy loss is exactly dominated by that of the Gaussian mechanism with noise multiplier $\sigma/\sqrt{T}$. Linear losses give an example such that DP-GD with last_iterate_only has exactly this privacy loss, since the final iterate reveals the sum of all the noisy gradient estimates. The worst-case privacy of DP-GD with intermediate_iterates is no worse than that of DP-GD with last_iterate_only. The reverse is also true (by postprocessing).
@@ Appendix D Implementation Details
For image classification model we used De et. al, [DBH+22] training approach with different batch sizes. We used Google Cloud A2-megagpu-16g machines with 16 Nvidia A100 40GB GPUs. Overall, we use roughly 60,000 GPU hours for our experiments. To compute the standard theoretical privacy bounds we use DP-accounting library  [^3] https://github.com/google/differential-privacy/tree/main. For each experiment we trained the model of image classification setting we trained 1000 models with and without the canary (i.e, differing example).
We implemented fine tuning with differential privacy on Gemma 2 [TMH+24] in JAX. The fine tuning is done in FP16. The attack here directly looks at the difference in the embedding table of the model for the tokens used as canaries before and after the finetuning. We trained 128 model with and without carries for language model fine tuning experiments.

### v3

@@ A.1 Computing $\delta$ from $\varepsilon$
These expressions can be evaluated efficiently. Since $f$ is monotone, it can be inverted via binary search. We can also use binary search to evaluate $\varepsilon$ as a function of $\delta$. Pseudocode for this is provided in the next subsection.
A.2 Pseudocode for computing $\varepsilon$ and $\delta$ values
[…]
[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines ]pythonanc/heuristic_code.py
@@ Appendix B Linear Worst Case for Full Batch Setting
We can formalize this using the language of Gaussian DP (Dong et al., 2019): DP-GD (Algorithm 1 with $q=1$) satisfies $\sqrt{T}/\sigma$-GDP. (Each iteration satisfies $1/\sigma$-GDP and adaptive composition implies the overall guarantee.) This means that the privacy loss is exactly dominated by that of the Gaussian mechanism with noise multiplier $\sigma/\sqrt{T}$. Linear losses give an example such that DP-GD with last_iterate_only has exactly this privacy loss, since the final iterate reveals the sum of all the noisy gradient estimates. The worst-case privacy of DP-GD with intermediate_iterates is no worse than that of DP-GD with last_iterate_only. The reverse is also true (by postprocessing).
@@ Appendix D Implementation Details
For image classification model we used De et. al, (De et al., 2022) training approach with different batch sizes. We used Google Cloud A2-megagpu-16g machines with 16 Nvidia A100 40GB GPUs. Overall, we use roughly 60,000 GPU hours for our experiments. To compute the standard theoretical privacy bounds we use DP-accounting library  [^3] https://github.com/google/differential-privacy/tree/main. For each experiment we trained the model of image classification setting we trained 1000 models with and without the canary (i.e, differing example).
We implemented fine tuning with differential privacy on Gemma 2 (Team et al., 2024) in JAX. The fine tuning is done in FP16. The attack here directly looks at the difference in the embedding table of the model for the tokens used as canaries before and after the finetuning. We trained 128 model with and without carries for language model fine tuning experiments.